Solve for z, and give your answer in the form atbi. 2+2 4i= 4x+11+19 2 = 0 matie
Answer #1

Georen expression Z+2-47 = 42+ 11 +19 Let z = atbi then, atbi+2-49 +2–49 – 4 Ca+bi) +1/+19i. (a+2) +66-4) (49+1) + (46+19) iA-4 a 1-2 -39 = 9 a= -3 a=-3 Also, -(6-4) -(6-4) = 4b+19. -btu 4b+19. -b-ub= 19-4. - 55 - 15 b = 15 -5 b -3 Hence, Z= a + b =so, z=-3-3i